If you get 5 Pineapples for $2.50, how many Pineapples you will get for $4.50 ?
Found 2 solutions by stanbon, Maths68:
Answer by stanbon(75887)

If you get 5 Pineapples for $2.50, how many Pineapples you will get for $4.50 ?
-----
x/450 = 5/250
------
x = 450(5/250)
------
x = 9 pineapples
========================
Cheers,
Stan H.

Answer by Maths68(1474)

In 2.5 we get 5 Pineapples
In 2.5/2.5 we get 5/2.5 Pineapples
In 1 we get 5/2.5 Pineapples
In 4.5*1 we get (4.5)(5/2.5) Pineapples
In 4.5 we get (45)(5/25) Pineapples
In 4.5 we get (45)(1/5) Pineapples
In 4.5 we get (45/5) Pineapples
In 4.5 we get 9 Pineapples